OpenStudy (anonymous):

Find the value of log2 1/32

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

how do you solve it? Can you show me the steps?

OpenStudy (anonymous):

first, i know that you know that log function has an inverse function of exponential function, you first convert it first in the form of exponential function to understand it better, or i say, it simpler in that way, log2 1/32=x convert to exponential form 2^x=1/32 2^x=1/2^5 2^x=2^-5 so x = -5 thats it. you can also solve it mentally when you master

OpenStudy (anonymous):

Thanks :)